# Transfer Learning Study of Motion Transformer-based Trajectory Predictions*

### v1

@@ II-A Related Work and Architectural Overview
[…]
MTR [1]
1st 2022
0.4129
0.6050
1.2207
0.1351
GTR-R36 [19]
3rd 2023
0.4255
0.6005
1.2225
0.1330
IAIR+ [18]
2nd 2023
0.4347
0.5783
1.1679
0.1238
MTR++ [14]
1st 2023
0.4329
0.5906
1.1939
0.1298
EDA [2]
newer
0.4401
0.5718
1.1702
0.1169
MGTR [3]
newer
0.4505
0.5918
1.2135
0.1298
@@ III-A General Setup
[…]
@@
Source / Target
Source
Target
Environment Setting
US Roads
German Roads
Number of Scenarios
575,205
190,933
Duration of Each Scenario
9 seconds
9 seconds
Trajectory Sampling Rate
10Hz
10Hz
Training Split
$84.\overline{6}$(487,005)
70% (133,653)
Validation Split
$7.\overline{6}$(44,100)
15% (28,640)
Test Split
$7.\overline{6}$(44,100)
15% (28,640)
Total Trajectories
2,566,096
190,933
Agent Split (Veh./Cyc./Ped.)
70/7/23%
100/0/0%
@@ IV-A Training & Evaluation Setup
CPU
Intel Core i7-13700K, 8x 3.40 GHz, 8xE, 30MB L3 Cache
MB
ASUS ROG STRIX Z790-F
GPU
NVIDIA RTX A5000, 24GB GDDR6
RAM
64GB (2x 32GB Kit) DDR5-5200 CL40
SSD
2TB Samsung 980 Pro, M.2 PCIe 4.0

### v2

# IEEE copyright notice

@@ II-A Related Work and Architectural Overview
MTR [1] 1st 2022 0.4129 0.6050 1.2207 0.1351
GTR-R36 [19] 3rd 2023 0.4255 0.6005 1.2225 0.1330
IAIR+ [18] 2nd 2023 0.4347 0.5783 1.1679 0.1238
MTR++ [14] 1st 2023 0.4329 0.5906 1.1939 0.1298
EDA [2] newer 0.4401 0.5718 1.1702 0.1169
MGTR [3] newer 0.4505 0.5918 1.2135 0.1298
@@ III-A General Setup
The study aims at investigating the knowledge transfer from the Waymo Open Motion Dataset (WOMD), refered to as source dataset $\textit{D}_{S}$ to the constructed CarMaker Dataset (CMD), refered to as target dataset $\textit{D}_{T}$. However, before the dataset specifics and the comparison in the Table II are discussed, the task is considered more formally in accordance with [29]. Hence, the datasets D represent observations of both a domain space and a task space, denoted as $\{(\mathcal{D},\mathcal{T})\}$. Here, the domain space $\mathcal{D}=\{\mathcal{X},P(X)\}$ encompasses an associated feature space $X$ and its corresponding marginal distribution $P(X)$, while the task space $\mathcal{T}=\{\mathcal{Y},f\}$ is defined through the label space $\mathcal{Y}$ and the implicit function $f:\mathcal{X}\rightarrow\mathcal{Y}$ to be learned.
Accordingly, the transfer learning objective can be stated as improving the learning of the implicit target function $f_{T}$ by using source observations $\textit{D}_{S}=\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}^{n_{S}}_{i=1}$ of $\{(\mathcal{D}_{S},\mathcal{T}_{S})\}$ and target observations $\textit{D}_{T}=\{(\boldsymbol{x}_{S},i,\boldsymbol{y}_{S},i)\}^{n_{T}}_{i=1}$ of $\{(\mathcal{D}_{T},\mathcal{T}_{T})\}$, with $\boldsymbol{x}_{j,i}\in\mathcal{X}_{j}\subseteq\mathbb{R}^{d_{X_{j}}}$, $\boldsymbol{y}_{j,i}\in\mathcal{Y}_{j}\subseteq\mathbb{R}^{d_{y_{j}}},j\in\{S,T\}$ and $n_{s},n_{T}\in\mathbb{N}^{+}$ where $0\ll n_{T}\ll n_{S}$. This highlights that unsupervised transfer learning ($n_{T}=0$) is not part of our study, as generating a small, specific dataset is not only possible but also meaningful [31] in the task at hand.
@@
Source / Target Source Target
Environment Setting US Roads German Roads
Number of Scenarios 575,205 190,933
Duration of Each Scenario 9 seconds 9 seconds
Trajectory Sampling Rate 10Hz 10Hz
Training Split $84.\overline{6}$(487,005) 70% (133,653)
Validation Split $7.\overline{6}$(44,100) 15% (28,640)
Test Split $7.\overline{6}$(44,100) 15% (28,640)
Total Trajectories 2,566,096 190,933
Agent Split (Veh./Cyc./Ped.) 70/7/23% 100/0/0%
@@ IV-A Training & Evaluation Setup
CPU Intel Core i7-13700K, 8x 3.40 GHz, 8xE, 30MB L3 Cache
MB ASUS ROG STRIX Z790-F
GPU NVIDIA RTX A5000, 24GB GDDR6
RAM 64GB (2x 32GB Kit) DDR5-5200 CL40
SSD 2TB Samsung 980 Pro, M.2 PCIe 4.0